# LRT-Diffusion: Calibrated Risk-Aware Guidance for Diffusion Policies

### v1

@@ Abstract
Keywords Offline Reinforcement learning; Diffusion Policies; Likelihood‑ratio Test; Inference; Risk‑aware Guidance; Calibration; Distribution Shift; Q‑guidance; Out‑of‑distribution Detection; D4RL.
1 Introduction
@@
Rather than always pulling samples toward the conditional head, we ask at every denoising step: is there enough evidence to move toward the “good‑action” direction for this state? Concretely, we split offline actions into a good subset and a background subset, train a two‑head diffusion model (an unconditional head on all data and a conditional head on good data) with class‑balancing and optional advantage‑based soft weights, and then gate the conditional pull at inference by a calibrated likelihood ratio. The gate is motivated by the Neyman–Pearson test NeymanPearson1933 : a single user knob Type‑I rate $\alpha$ controls the tolerated false activations under $H_{0}$ (“background is correct”); while the hard LRT is UMP at level $\alpha$ under equal covariances, we use a smooth gate in practice for numerical stability and keep $\alpha$ interpretable via calibration on held‑out states.
Method overview.
@@
We introduce LRT‑Diffusion, a risk‑aware, inference‑only sampling scheme for diffusion policies. At each denoising step we make a binary decision between a background prediction and a data‑conditioned prediction. We accumulate evidence during sampling and open the gate only when the evidence is strong, interpolating between the two predictions with a data‑dependent weight. A single threshold is calibrated once so that the empirical false‑activation rate does not exceed a user‑chosen $\alpha$, turning guidance from a fixed push into an evidence‑driven adjustment: weak evidence $\Rightarrow$ stay near the background prior; strong evidence $\Rightarrow$ move decisively toward the conditional policy. Training is unchanged.
Composition with value guidance.
@@
We optionally combine LRT gating with a small critic step: at each denoising update, take a capped step that increases the learned critic at the current proposal (standard decreasing schedule). The gradient can be evaluated near the background proposal, the LRT-gated proposal, or an interpolation; the threshold is calibrated with the same choice, so the Type-I semantics are preserved.
Theoretical guarantees at a glance.
@@
Under equal‑covariance heads, the hard gate(thresholding under LRT assumptions) is the uniformly most powerful test at level $\alpha$; our soft gate (a numerically stable surrogate) retains the threshold’s semantics and improves stability. Beyond testing optimality, we analyze offline‑RL error propagation and show that reducing the policy’s state‑conditional OOD rate—which LRT lowers by construction—tightens a lower bound on true return relative to pure QQ‑guidance when the critic is unreliable off‑support.
Practical picture.
@@
The gate exposes monotone, interpretable knobs: $\alpha$ (risk), $\beta_{\max}$ (max pull), and $\delta$ (sharpness). LRT can be combined with a small critic step evaluated at the background mean, the LRT-gated mean, or a blend; the threshold is calibrated with that same choice so the Type-I semantics are preserved. Empirically on D4RL MuJoCo, LRT tracks the target $\alpha$ and improves the return–OOD Pareto relative to strong $Q$-guided baselines in our implementation.
1.1 Contributions
Calibrated LRT guidance for diffusion policies.
@@ 1.1 Contributions
Unlike prior research integrating value or energy guidance into the diffusion model, we cast each denoising step as a likelihood-ratio test between a background head and a “good” head, and calibrate a single threshold $\tau$ on held-out states to bound the trajectory-level Type-I rate at a user-chosen $\alpha$. Training remains vanilla (two heads with class balancing and optional advantage weights).
Labeling-and-weighting recipe for the conditional head.
@@
Rather than fitting a single conditional diffusion model to all (state, action) pairs, as is common in return-dominated diffusion or $Q$-guided training, we split offline actions into a good subset and a background subset using IQL advantages (top-$p$ quantile), and train a two-head diffusion model with class-balancing and advantage-aware soft weights. This improves the quality of the conditional head without extra supervision.  [^1] E.g., Diffuser/Decision‑Diffuser use return conditioning and sampling-time guidance; Diffusion‑QL injects value signals into the training loss, all without a binary good/background head.
Actionable theory with finite-sample and stability guarantees.
@@
Under equal covariances the hard LRT is UMP at level $\alpha$ (Prop. 1); a one-shot Monte-Carlo calibration yields $\mathbb{P}_{H_{0}}(\ell_{\mathrm{cum}}\geq\hat{\tau})\leq\alpha+\varepsilon_{n}$ (Thm. 2). We bound the per-step deterministic drift and show the cumulative LLR is sub-Gaussian with an explicit variance proxy (Sec. 5.2), explaining why calibration is numerically stable.
Return bounds that connect risk control to distribution shift.
@@
With a standard offline error split, we prove a comparison bound for true returns between LRT-guided and $Q$-guided sampling (Prop. 3). A mild monotonicity assumption turns the calibrated level $\alpha$ into a conservative OOD upper bound (Prop. 4), clarifying when and why LRT is preferable to pure $Q$-guidance.
Compatibility with value gradients via matched calibration.
@@
LRT controls risk while the QQ‑step pursues return, and these roles are orthogonal. We show that composing LRT with a small $Q$-step preserves level-$\alpha$ semantics so long as calibration uses the exact deployed sampler (Prop. 2). Thus, risk control (through $\alpha$ and its calibrated $\tau$) is decoupled from return‑seeking (through $\nabla_{a}\hat{Q}$), making the QQ‑update a plug‑and‑play module. The knobs $(\alpha,\beta_{\max},\delta)$ remain monotone and interpretable.
Empirical validation.
@@
On D4RL MuJoCo, LRT-Diffusion honors the target Type-I rate and improves the return–OOD Pareto over strong $Q$-guided baselines in our implementation, with ablations over $\alpha$, $\beta_{\max}$, $\delta$, and gradient anchoring.
2 Related Work
Distribution shift in offline RL and conservative learning.
@@ 2 Related Work
A central obstacle in offline RL is that policy actions may fall outside the behavior support, where critic estimates are brittle. Conservative algorithms explicitly counter this by penalizing or constraining OOD actions, e.g., BCQ Fujimoto2019BCQ , BEAR Kumar2019BEAR , CQL Kumar2020CQL , and TD3+BC Fujimoto2021TD3BC . IQL Kostrikov2022IQL learns values via expectile regression and performs advantage-weighted improvement without explicit behavior cloning. Our work is complementary: we keep training unchanged (e.g., use IQL advantages only to label/weight “good” actions) and control OOD at sampling time via a calibrated gate.
Diffusion policies for decision-making and their guidance.
@@
Diffusion models have been adopted for control and planning (e.g., Diffuser Janner2022Diffuser ). To improve returns, many systems apply Q-guidance—adding an action-space step along $\nabla_{a}\hat{Q}(s,a)$ with hand-tuned schedules and clips—conceptually analogous to guidance in image diffusion (e.g., classifier-based guidance Dhariwal2021GuidedDiffusion and classifier-free guidance Ho2022CFG ). However, such heuristics lack a statistical notion of risk: there is no global control on the probability of “falsely” pulling away from the background prior.
Risk-aware control and calibration.
@@
Risk-sensitive and conservative offline RL methods (e.g., behavior-regularized or pessimistic objectives) encode risk in the training loss by penalizing value estimates or constraining policy deviation from the dataset (Fujimoto2019BCQ, ; Wu2019BRAC, ; Kumar2020CQL, ; Nair2020AWAC, ; Peng2019AWR, ; Kostrikov2022IQL, ). A complementary line uses distribution-free calibration to turn data-driven thresholds into finite-sample guarantees (e.g., conformal prediction) (Vovk2005ALRW, ; Angelopoulos2023ConformalSurvey, ). In diffusion-based decision making, sampling is typically guided heuristically (e.g., $Q$-gradient pushes with hand-tuned schedules/clipping) without an explicit statistical notion of risk. Our approach bridges these threads: we cast each reverse step as a simple-vs-simple test between an unconditional (background) head and a conditional (good) head, accumulate a log-likelihood ratio (LLR), and calibrate one threshold on held-out states so that the empirical Type-I rate under $H_{0}$ does not exceed a user-chosen $\alpha$. Under equal covariances—satisfied by our two-head design—the hard likelihood-ratio test is uniformly most powerful (Neyman–Pearson) (NeymanPearson1933, ), and we use a smooth gate for numerical stability while preserving the same $\alpha$-semantics via calibration. This gives a statistically grounded alternative to heuristic mixing: guidance becomes evidence-driven with an explicit level-$\alpha$ risk budget at inference time, without changing training.
Hypothesis testing and likelihood ratios in RL
@@
Statistical testing has been used in RL primarily for off-policy evaluation and safe policy improvement, not for stepwise guidance inside a generative sampler. High-confidence OPE Thomas2015HCOPE derives concentration-based bounds to accept/reject policy deployment with user-specified risk, while SPIBB Laroche2019SPIBB ensures safe improvement over a baseline via count-based constraints rather than explicit tests. In safety-constrained RL, trust-region and constraint-violation controls (e.g., CPO Achiam2017CPO ) target reliable deployment but do not implement likelihood-ratio gating during action generation. Classical sequential tests such as Wald’s SPRT Wald1945SPRT have inspired sequential decision criteria, yet we are not aware of prior work that uses a calibrated likelihood-ratio test as an inference-time gate to steer diffusion-policy denoising steps. Our method fills this gap by turning denoising guidance into an evidence-controlled, level-$\alpha$ process.
Advantage-based labeling and weighting.
@@
For higher return actions, advantage-weighted and critic-regularized schemes (RWR Peters2007RWR , AWR Peng2019AWR , AWAC Nair2020AWAC ) are proposed. We use IQL advantages to (i) define a top-$p$ “good” subset for the conditional head and (ii) optionally apply a temperature-controlled soft weight on positives, while keeping a background head trained on all data. This two-head setup strengthens the conditional signal without altering the base training pipeline, and remains compatible with our calibrated LRT at inference. Our two-head design enforces equal covariances by construction, which simplifies the LLR and aligns with the NP test.
3 Background
@@
Offline RL is given a fixed dataset $\mathcal{D}=\{(s^{(i)},a^{(i)},r^{(i)},s^{(i)^{\prime}})\}_{i=1}^{N}$, collected by an unknown behavior policy $\pi$, and aims to learn a high-return policy while controlling out-of-distribution risk. In the dataset, $s\in\mathbb{R}^{d_{s}}$ is the current state, $a\in\mathbb{R}^{d_{a}}$ is the action, $r\in\mathbb{R}$ is the reward, $s^{\prime}$ is the next state. We consider the process as a discounted Markov Decision Process with continuous state/action spaces.
3.1 IQL Advantages and “Good vs. Background” Labels
@@
For each pair of state and action $(s,a)\in\mathcal{D}$, we train an IQL-style critic Kostrikov2022IQL $(\hat{Q}(s,a),\hat{V}(s))$ on standardized inputs and define an advantage $A(s,a)$ as:
$A(s,a)\;=\;\hat{Q}(s,a)-\hat{V}(s),$
which quantifies the value contributed by action $a$ at state $s$ relative to the state’s baseline.
Let $\kappa$ be the (1p)(1-p)-quantile of $\{A(s_{i},a_{i})\}_{i=1}^{N}$. We label each pair of state-action data by
$c\;=\;\mathbf{1}\{A(s,a)\geq\kappa\}\in\{0,1\}.$
@@
Pairs with c=1c{=}1 are considered good, as their advantages fall within the top (1p)(1-p) fraction of the dataset, while those with c=0c{=}0 form the non-advantage subset. In practice, we set $p=0.2$, allocating approximately $20\%$ of the data as good examples.
3.2 Diffusion Policies with Two Heads
@@
We train a diffusion policy with two heads: an unconditional head trained on all state–action pairs $\mathcal{D}$, and a conditional head trained only on the high-advantage subset (c=1c{=}1) we labeled before. Formally, let $\mathcal{D}_{\mathrm{all}}:=\mathcal{D}$ and $\mathcal{D}_{\mathrm{good}}:=\{(s,a,r,s^{\prime})\in\mathcal{D}:c=1\}$. The unconditional head learns broad dataset coverage from $\mathcal{D}_{\mathrm{all}}$, while the conditional head specializes to high-advantage behavior using $\mathcal{D}_{\mathrm{good}}$.
3.2.1 Two heads and notation
@@
We use a shared backbone $\phi_{\theta}(s,a_{t},t)$ with two output branches (“heads”): an unconditional/background head trained on $\mathcal{D}_{\mathrm{all}}$, and a conditional/good head trained on $\mathcal{D}_{\mathrm{good}}$. Each head maps features to a DDPM-style noise prediction,
@@
3.2.2 DDPM parameterization
We adopt the predict-$\varepsilon$ parameterization Ho2020DDPM ; Nichol2021ImprovedDDPM with schedule $\{\alpha_{t},\bar{\alpha}_{t}\}_{t=1}^{T}$, where $\bar{\alpha}_{t}=\prod_{\tau=1}^{t}\alpha_{\tau}$. Given the current latent action $a_{t}$ at step $t$, a noise prediction $\hat{\varepsilon}$ induces a Gaussian reverse kernel with shared covariance $\sigma_{t}^{2}I$ and mean
@@ 3.2.2 DDPM parameterization
We define the head-specific means
@@
$a_{t-1}\mid a_{t},s~\sim~\mathcal{N}\!\big(\mu_{u}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),\quad a_{t-1}\mid a_{t},s~\sim~\mathcal{N}\!\big(\mu_{c}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$
@@
Notation clarification.
We write dataset indices as parenthesized superscripts, e.g., $a^{(i)}$ is the ithi-th action in the dataset $\mathcal{D}$. We reserve subscripts $t$ for diffusion steps, e.g., $a_{t}$ is the latent at step $t$ and $a_{0}$ is the final action..
3.3 Hypotheses for a Reverse Step

### v2

@@ Abstract.
Key words and phrases:
Offline Reinforcement learning; Diffusion Policies; Likelihood‑ratio Test; Inference; Risk‑aware Guidance; Calibration; Distribution Shift; Q‑guidance; Out‑of‑distribution Detection; D4RL.
doi:
VBFF4869
This paper has been accepted to AAMAS 2026.
1. Introduction
@@
Intuition. Rather than always pulling samples toward the conditional head, we ask at every denoising step: is there enough evidence to move toward the “good‑action” direction for this state? Concretely, we split offline actions into a good subset and a background subset, train a two‑head diffusion model (an unconditional head on all data and a conditional head on good data) with class‑balancing and optional advantage‑based soft weights, and then gate the conditional pull at inference by a calibrated likelihood ratio. The gate is motivated by the Neyman–Pearson test NeymanPearson1933: a single user knob Type‑I rate $\alpha$ controls the tolerated false activations under $H_{0}$ (“background is correct”); while the hard LRT is UMP at level $\alpha$ under equal covariances, we use a smooth gate in practice for numerical stability and keep $\alpha$ interpretable via calibration on held‑out states.
@@
Method overview. We introduce LRT‑Diffusion, a risk‑aware, inference‑only sampling scheme for diffusion policies (see Fig 1). At each denoising step we make a binary decision between a background prediction and a data‑conditioned prediction. We accumulate evidence during sampling and open the gate only when the evidence is strong, interpolating between the two predictions with a data‑dependent weight. A single threshold is calibrated once per task so that the empirical false‑activation rate does not exceed a user‑chosen $\alpha$, turning guidance from a fixed push into an evidence‑driven adjustment: weak evidence $\Rightarrow$ stay near the background prior; strong evidence $\Rightarrow$ move decisively toward the conditional policy. Training is unchanged. A detailed, step-by-step illustration of the inference-time pipeline is provided in Fig. 4.
Figure 1. Overall pipeline of LRT-Diffusion. Training is vanilla; all risk control is applied at inference via a calibrated gate.
@@
Composition with value guidance. We optionally combine LRT gating with a small critic step: at each denoising update, take a capped step that increases the learned critic at the current proposal (standard decreasing schedule). The gradient can be evaluated near the background proposal, the LRT-gated proposal, or an interpolation; the threshold is calibrated with the same choice, so the Type-I semantics are preserved.
@@
Theoretical guarantees at a glance. Under equal‑covariance heads, the hard gate is the uniformly most powerful test at level $\alpha$; our soft gate (a numerically stable surrogate) retains the threshold’s semantics and improves stability. Beyond testing optimality, we analyze offline‑RL error propagation and show that reducing the policy’s state‑conditional OOD rate—which LRT lowers in many regimes—tightens a lower bound on true return relative to pure QQ‑guidance when the critic is unreliable off‑support.
@@
Practical picture. The gate exposes monotone, interpretable knobs: $\alpha$ (risk), $\beta_{\max}$ (max pull), and $\delta$ (sharpness). LRT can be combined with a small critic step evaluated at the background mean, the LRT-gated mean, or a blend; the threshold is calibrated with that same choice so the Type-I semantics are preserved. Empirically on D4RL MuJoCo, LRT tracks the target $\alpha$ and yields a calibrated return-risk frontier: LRT serves as a low-risk anchor, while LRT+Q increases return along the frontier.
1.1. Contributions
@@
Calibrated LRT guidance for diffusion policies. Unlike prior research integrating value or energy guidance into the diffusion model, we cast each denoising step as a likelihood-ratio test between a background head and a “good” head, and calibrate a single threshold $\tau$ on held-out states to bound the trajectory-level Type-I rate at a user-chosen $\alpha$. Training remains vanilla (two heads with class balancing and optional advantage weights).
@@
Labeling-and-weighting recipe for the conditional head. Rather than fitting a single conditional diffusion model to all (state, action) pairs, as is common in return-dominated diffusion or $Q$-guided training, we split offline actions into a good subset and a background subset using IQL advantages (top-$p$ quantile), and train a two-head diffusion model with class-balancing and advantage-aware soft weights. This improves the quality of the conditional head without extra supervision.  [^1] E.g., Diffuser/Decision‑Diffuser use return conditioning and sampling-time guidance; Diffusion‑QL injects value signals into the training loss, all without a binary good/background head.
@@
Actionable theory with finite-sample and stability guarantees. Under equal covariances the hard LRT is UMP at level $\alpha$ (Prop. 5.1); a one-shot Monte-Carlo calibration yields $\mathbb{P}_{H_{0}}(\ell_{\mathrm{cum}}\geq\hat{\tau})\leq\alpha+\varepsilon_{n}$ (Thm. 5.2.1). We bound the per-step deterministic drift and show the cumulative LLR is sub-Gaussian with an explicit variance proxy (Sec. 5.2), explaining why calibration is numerically stable.
@@
Return bounds that connect risk control to distribution shift. With a standard offline error split, we prove a comparison bound for true returns between LRT-guided and $Q$-guided sampling (Prop. 5.3). A mild monotonicity assumption turns the calibrated level $\alpha$ into a conservative OOD upper bound (Prop. 5.3), clarifying when and why LRT is preferable to pure $Q$-guidance.
@@
Compatibility with value gradients via matched calibration. LRT controls risk while the QQ‑step pursues return, and these roles are orthogonal. We show that composing LRT with a small $Q$-step preserves level-$\alpha$ semantics so long as calibration uses the exact deployed sampler (Prop. 5.1). Thus, risk control (through $\alpha$ and its calibrated $\tau$) is decoupled from return‑seeking (through $\nabla_{a}\hat{Q}$), making the QQ‑update a plug‑and‑play module. The knobs $(\alpha,\beta_{\max},\delta)$ remain monotone and interpretable.
@@
Empirical validation. On D4RL MuJoCo, LRT-Diffusion honors the target Type-I rate and yields a calibrated return-risk frontier, with ablations over $\alpha$, $\beta_{\max}$, $\delta$, and gradient anchoring.
2. Related Work
@@
Distribution shift in offline RL and conservative learning. A central obstacle in offline RL is that policy actions may fall outside the behavior support, where critic estimates are brittle. Conservative algorithms explicitly counter this by penalizing or constraining OOD actions, e.g. (Fujimoto2019BCQ; Kumar2019BEAR; Kumar2020CQL; Fujimoto2021TD3BC; Kostrikov2022IQL) learns values via expectile regression and performs advantage-weighted improvement without explicit behavior cloning. Our work is complementary: we keep training unchanged (e.g., use IQL advantages only to label/weight “good” actions) and control OOD at sampling time via a calibrated gate.
@@
Diffusion policies for decision-making and their guidance. Diffusion models have been adopted for control and planning Janner2022Diffuser. To improve returns, many systems apply Q-guidance—adding an action-space step along $\nabla_{a}\hat{Q}(s,a)$ with hand-tuned schedules and clips—conceptually analogous to guidance in image diffusion Dhariwal2021GuidedDiffusion; Ho2022CFG. However, such heuristics lack a statistical notion of risk: there is no global control on the probability of “falsely” pulling away from the background prior.
@@
Risk-aware control and calibration. Risk-sensitive and conservative offline RL methods (e.g., behavior-regularized or pessimistic objectives) encode risk in the training loss by penalizing value estimates or constraining policy deviation from the dataset (Fujimoto2019BCQ; Wu2019BRAC; Kumar2020CQL; Nair2020AWAC; Peng2019AWR; Kostrikov2022IQL). A complementary line uses distribution-free calibration to turn data-driven thresholds into finite-sample guarantees (e.g., conformal prediction) (Vovk2005ALRW; Angelopoulos2023ConformalSurvey). In diffusion-based decision making, sampling is typically guided heuristically (e.g., $Q$-gradient pushes with hand-tuned schedules/clipping) without an explicit statistical notion of risk. Our approach bridges these threads: we cast each reverse step as a simple-vs-simple test between an unconditional (background) head and a conditional (good) head, accumulate a log-likelihood ratio (LLR), and calibrate one threshold on held-out states so that the empirical Type-I rate under $H_{0}$ does not exceed a user-chosen $\alpha$. Under equal covariances—satisfied by our two-head design—the hard likelihood-ratio test is uniformly most powerful (NeymanPearson1933), and we use a smooth gate for numerical stability while preserving the same $\alpha$-semantics via calibration. This gives a statistically grounded alternative to heuristic mixing: guidance becomes evidence-driven with an explicit level-$\alpha$ risk budget at inference time, without changing training.
@@
Hypothesis testing and likelihood ratios in RL Testing in RL is typically used for deployment-time decisions (high-confidence OPE Thomas2015HCOPE, safe improvement such as SPIBB Laroche2019SPIBB), rather than within a generative sampler. Safety-constrained methods (e.g., CPO Achiam2017CPO) control violations but do not apply likelihood-ratio gating during action generation. Although sequential tests like SPRT Wald1945SPRT inspire sequential criteria, we are not aware of calibrated likelihood-ratio gates used to steer diffusion denoising at inference. Our approach turns guidance into an evidence-controlled, level-$\alpha$ procedure.
@@
Advantage-based labeling and weighting. For higher return actions, advantage-weighted and critic-regularized schemes ( (Peters2007RWR; Peng2019AWR; Nair2020AWAC)) are proposed. We use IQL advantages to (i) define a top-$p$ “good” subset for the conditional head and (ii) optionally apply a temperature-controlled soft weight on positives, while keeping a background head trained on all data. This two-head setup strengthens the conditional signal without altering the base training pipeline, and remains compatible with our calibrated LRT at inference. Our two-head design enforces equal covariances by construction, which simplifies the LLR and aligns with the NP test.
3. Background
@@
Offline RL is given a fixed dataset $\mathcal{D}=\{(s^{(i)},a^{(i)},r^{(i)},s^{(i)^{\prime}})\}_{i=1}^{N}$, collected by an unknown behavior policy $\pi$, and aims to learn a high-return policy while controlling out-of-distribution risk. In the dataset, $s\in\mathbb{R}^{d_{s}}$ is the current state, $a\in\mathbb{R}^{d_{a}}$ is the action, $r\in\mathbb{R}$ is the reward, $s^{\prime}$ is the next state. We treat the process as a discounted Markov Decision Process with continuous state/action spaces.
3.1. IQL Advantages and “Good vs. Background” Labels
@@
For each pair of state and action $(s,a)\in\mathcal{D}$, we train an IQL-style critic Kostrikov2022IQL $(\hat{Q}(s,a),\hat{V}(s))$ on standardized inputs and define an advantage $A(s,a)$ as $A(s,a)\;=\;\hat{Q}(s,a)-\hat{V}(s)$, which quantifies the value contributed by action $a$ at state $s$ relative to the state’s baseline. Let $\kappa$ be the top $(p)$-quantile of $\{A(s_{i},a_{i})\}_{i=1}^{N}$. We label each pair of state-action data by $c\;=\;\mathbf{1}\{A(s,a)\geq\kappa\}\in\{0,1\}$.
@@
Pairs with c=1c{=}1 are considered good, as their advantages fall within the top $(p)$ fraction of the dataset, while those with c=0c{=}0 form the non-advantage subset. In practice, we set $p=0.2$, allocating approximately $20\%$ of the data as good examples.
3.2. Diffusion Policies with Two Heads
@@
We train a diffusion policy with two heads: an unconditional head trained on all state–action pairs $\mathcal{D}$, and a conditional head trained only on the high-advantage subset we labeled beforehand. Formally, let $\mathcal{D}_{\mathrm{all}}:=\mathcal{D}$ and $\mathcal{D}_{\mathrm{good}}:=\{(s,a,r,s^{\prime})\in\mathcal{D}:c=1\}$. The unconditional head learns broad dataset coverage from $\mathcal{D}_{\mathrm{all}}$, while the conditional head specializes to advantage behavior using $\mathcal{D}_{\mathrm{good}}$.
@@
Two heads and notation. We use a shared backbone $\phi_{\theta}(s,a_{t},t)$ with two output branches (“heads”): an unconditional/background head trained on $\mathcal{D}_{\mathrm{all}}$, and a conditional/good head trained on $\mathcal{D}_{\mathrm{good}}$. Each head maps features to a DDPM-style noise prediction,
@@
DDPM parameterization. We adopt the predict-$\varepsilon$ parameterization Ho2020DDPM; Nichol2021ImprovedDDPM with schedule $\{\alpha_{t},\bar{\alpha}_{t}\}_{t=1}^{T}$, where $\bar{\alpha}_{t}=\prod_{\tau=1}^{t}\alpha_{\tau}$. Given the current latent action $a_{t}$ at step $t$, a noise prediction $\hat{\varepsilon}$ induces a Gaussian reverse kernel with shared covariance $\sigma_{t}^{2}I$ and mean
@@
Throughout this subsection, we assume that the two diffusion heads share the same reverse-time variance $\sigma_{t}^{2}$, as given by the diffusion schedule, which enables a closed-form expression for the step-wise log-likelihood ratio. We discuss this equal-variance choice and examine robustness when it is deliberately violated at inference time in Sec. 6.5. We define the head-specific means
@@
$\tiny a_{t-1}\mid a_{t},s\penalty 10000\ \sim\penalty 10000\ \mathcal{N}\!\big(\mu_{u}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),\quad a_{t-1}\mid a_{t},s\penalty 10000\ \sim\penalty 10000\ \mathcal{N}\!\big(\mu_{c}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$
@@
Notation clarification. We write dataset indices as parenthesized superscripts, e.g., $a^{(i)}$ is the ithi-th action in the dataset $\mathcal{D}$. We reserve subscripts $t$ for diffusion steps, e.g., $a_{t}$ is the latent at step $t$ and $a_{0}$ is the final action..
3.3. Hypotheses for a Reverse Step